bob makes 20 drinks in five minutes, sarah makes 20 drinks in 10 minutes, and Gary makes 20 drinks in 15 minutes. How long will it take them to make the 20 drinks if they were working together?
Answer by edjones(4080)

bob does 1/5 of the job in 1 minute.
sarah does 1/10 otj in 1 min.
gary does 1/15 otj in 1 min.
.
1/5+1/10+1/15=6/30+3/30+2/30=11/30 otj is done by the three of the in 1 minute.
So it takes 30/11 min to complete the job or 2 8/11 min or about 2 min 44 sec.